OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = 4x – 5, what is f(x) ◦ g(x)? a. 18x + 3 b. 8x – 7 c. 8x – 15 d. 8x + 7

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow (2x + 3)(4x - 5) }\) \(\Large \color{purple}{\rightarrow 2x(4x - 5) + 3(4x - 5) }\) Solve now?

OpenStudy (anonymous):

yes i ddi FoIL and got a werid answer...

OpenStudy (anonymous):

8x - 10x + 12x - 15

Parth (parthkohli):

Let me check it: \(\Large \color{Midnightblue}{\rightarrow 8x^2 - 10x + 12x - 15 }\) \(\Large \color{Midnightblue}{\rightarrow 8x^2 + 2x - 15 }\)

OpenStudy (anonymous):

but thats what i got and taht is none of the choice/...

Parth (parthkohli):

Hmm :/

Parth (parthkohli):

@lgbasallote @inkyvoyd

OpenStudy (anonymous):

no... actually, it is like this... \[f(x).g(x)=f(g(x))\] \[=f(4x-5)=2(4x-5)+3\] now, try solve this...

OpenStudy (anonymous):

oh i get it so its 8x - 10 + 3 = 8x - 7

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

yes, correct...=)

OpenStudy (anonymous):

tahnk u!

OpenStudy (anonymous):

you are welcome...=)